Use the product rule to simplify the expression. 275 25.VII 276 5v 55
Answer #1

Solution:-

Given that,

\sqrt{275}

Let's simplify the expression -

\sqrt{275}

Prime\; \; factorization\; \; of\; \; 275:\; 5^2\times 11\; \; \; \; (i.e\; 5\times 5\times 11=275)

Therefore,

\sqrt{275}

\Rightarrow \sqrt{5^2\times 11}

Applying radical rule:

  • \sqrt[n]{ab} = \sqrt[n]{a}\times \sqrt[n]{b}

Therefore,

\sqrt{5^2\times 11}

\Rightarrow \sqrt{5^2}\times \sqrt{11}​​​​​​​

Applying radical rule:

  • D=​​​​​​​

Therefore,

\sqrt{5^2}\times \sqrt{11}

\Rightarrow \boldsymbol{5\sqrt{11}}

Hence,

\sqrt{275}=\boldsymbol{5\sqrt{11}}\; \; \; \; \; \; \; \; (Answer)